### v1

# Can Watermarked LLMs be Identified by Users via Crafted Prompts?

Abstract
Text watermarking for Large Language Models (LLMs) has made significant progress in detecting LLM outputs and preventing misuse. Current watermarking techniques offer high detectability, minimal impact on text quality, and robustness to text editing. However, current researches lack investigation into the imperceptibility of watermarking techniques in LLM services. This is crucial as LLM providers may not want to disclose the presence of watermarks in real-world scenarios, as it could reduce user willingness to use the service and make watermarks more vulnerable to attacks. This work is the first to investigate the imperceptibility of watermarked LLMs. We design an identification algorithm called Water-Probe that detects watermarks through well-designed prompts to the LLM. Our key motivation is that current watermarked LLMs expose consistent biases under the same watermark key, resulting in similar differences across prompts under different watermark keys. Experiments show that almost all mainstream watermarking algorithms are easily identified with our well-designed prompts, while Water-Probe demonstrates a minimal false positive rate for non-watermarked LLMs. Finally, we propose that the key to enhancing the imperceptibility of watermarked LLMs is to increase the randomness of watermark key selection. Based on this, we introduce the Water-Bag strategy, which significantly improves watermark imperceptibility by merging multiple watermark keys.
1 Introduction
The rapid advancement of large language models (LLMs) has led to remarkable achievements in tasks such as question answering (Zhuang et al., 2024), programming (Jiang et al., 2024b), and reasoning (Wei et al., 2022), with widespread applications across various scenarios. However, the extensive use of LLMs has also raised concerns regarding copyright protection and misuse. Recent research indicates that malicious attackers can steal LLMs through model extraction techniques (Yao et al., 2024), and some users may abuse LLMs to generate and spread harmful information (Wei et al., 2024).
Text watermarking techniques for LLMs have become an important method to mitigate the above issues by adding detectable features to LLM outputs (Liu et al., 2024c). Recent researches on LLM watermarking have focused on improving watermark detectability (Kirchenbauer et al., 2023a), minimizing impact on generated text (Aaronson & Kirchner, 2022), and enhancing robustness against text modifications (Liu et al., 2024b). However, no work has considered the imperceptibility of watermarked LLMs, i.e., whether users can know if an LLM service is watermarked. In real-world scenarios, LLM service providers may not disclose the existence of watermarks, as it could reduce user willingness to use the service and make it more vulnerable to attacks (Sadasivan et al., 2023). As more LLM services consider implementing watermarks, it is crucial to investigate whether users can identify watermarked LLMs solely through crafted prompts.
Some studies focus on the imperceptibility of watermarked text, ensuring watermarked and non-watermarked texts are indistinguishable (Hu et al., 2023; Wu et al., 2023b). However, even if individual watermarked texts are imperceptible, the distribution of numerous watermarked texts may reveal whether the LLM is watermarked, especially when repeatedly sampling with the same watermark key (Wu et al., 2024). While some studies explore cracking watermarks using large volumes of watermarked text (Jovanović et al., 2024; Sadasivan et al., 2023; Wu & Chandrasekaran, 2024; Zhang et al., 2024b), they assume the target LLM is watermarked and cannot determine if the LLM is watermarked. Research on the imperceptibility of watermarked LLMs should investigate whether watermarked and non-watermarked LLMs can be distinguished.
Figure 1: Illustration of our Water-Probe algorithm for identifying watermarked LLMs. We first construct two prompts with similar output distributions, then sample repeatedly using two fixed watermark keys for each prompt. The presence of a watermark is determined by comparing the similarity of distribution differences between the two prompts. Details in section 4.
In this work, we conduct the first study on the imperceptibility of watermarked LLMs. We propose the Water-Probe algorithm to identify watermarked LLMs under two main types of watermarking methods. Our motivation stems from a key observation: current LLM watermarking algorithms expose consistent bias when repeatedly sampled under the same watermark key. Based on this, our Water-Probe algorithm first crafts prompts to perform repeated sampling under the same watermark key, then compares the consistency of sampling distribution differences across different prompts under a pair of watermark keys. Highly consistent differences indicate a watermarked LLM.
In our experiments, we demonstrate that the Water-Probe algorithm achieves high accuracy in detecting various types of watermarked LLMs. We also show its applicability across different LLMs, maintaining a low false positive rate for non-watermarked LLMs. Furthermore, our algorithm exhibits robust performance across different sampling methods and temperature settings.
Finally, we explore methods to enhance the imperceptibility of watermarked LLMs. We find that increasing the randomness of watermark key selection is crucial, as it makes it more difficult to construct prompts for repeated sampling using the same key. Based on this, we propose the Water-Bag algorithm, which combines multiple watermark keys into one, randomly selecting a key for each generation and choosing the highest score for detection. However, increasing key selection randomness often leads to a slight decrease in detectability. Addressing this trade-off should be an important direction for future work.
Our main contributions can be summarized as follows:
  • We conduct the first study on the imperceptibility of watermarked LLMs, proposing the Water-Probe algorithm to effectively identify watermarked LLMs by exploiting consistent biases in watermarking keys.
  • We demonstrate the high accuracy and robustness of our algorithm across various LLMs, watermarking methods, and generation settings, while maintaining a minimal false positive rate for non-watermarked LLMs.
  • We introduce the Water-Bag algorithm to enhance LLM watermark imperceptibility.
2 Related Work
Large language model (LLM) watermarking techniques (Liu et al., 2024c) have become crucial for copyright protection (Sander et al., 2024), generated text detection (Wu et al., 2023a), and preventing misuse (Liu et al., 2024d). These watermarking algorithms introduce subtle features during text generation that are almost imperceptible to humans but detectable by designed detection algorithms. LLM watermarking can be broadly categorized into two types. The first type is post-processing watermarking, which modifies generated text using format-based (Sato et al., 2023), lexical-based (Yang et al., 2023), syntax-based (Wei et al., 2022), or generation-based (Zhang et al., 2024a) approaches to add watermark. However, post-processing watermarking methods require waiting for text generation to complete before modifying and adding watermarks, which is not suitable for current LLM services that require real-time text generation. Another category of watermarking algorithms, known as watermarking during generation, typically involves adjusting the distribution of the next generated token based on a watermark key. For instance, the KGW (Kirchenbauer et al., 2023a) algorithm divides the vocabulary into red and green lists, increasing the probability of tokens in the green list. SIR (Liu et al., 2024b) further modifies logits based on semantic information, enhancing watermark robustness. UPV (Liu et al., 2024a) utilizes two independent networks for watermark embedding and detection, enabling public verification. Additionally, some research has explored watermarking techniques for code generation scenarios (Lu et al., 2024) or perform watermark detection on longer documents (Pan et al., 2024b) or even translated documents (He et al., 2024). An important objective of these methods is to maintain the imperceptibility of generated text, i.e., watermarked and non-watermarked text should have identical distributions, with some distortion-free algorithms showing promising results (Kuditipudi et al., 2023; Aaronson & Kirchner, 2022). However, previous work has overlooked the imperceptibility of watermarked LLMs themselves, i.e., whether external users can detect if an LLM service contains watermarks without disclosure. This work investigates the imperceptibility of watermarked LLMs.
3 Preliminaries
To facilitate subsequent discussion, we introduce LLMs and LLM watermarking algorithms, then present the objectives of this work. We first provide the definition of LLM used in this work as follows:
Definition 1 (Large Language Model).
An LLM $M$ is a function that, given an input $x$ and a partial output sequence $y_{1:i-1}$, produces a probability distribution $P_{M}(y_{i}|x,y_{1:i-1})$ over possible next tokens $y_{i}$. The model generates complete outputs by iteratively sampling from these distributions.
Since this work focuses solely on watermarks embedded in LLM services, we will only consider the watermark during generation type mentioned in section 2. The definition of the watermark rule is given below.
Definition 2 (Watermark Rule).
A watermark rule is typically a function $F$ that adjusts the current LLM’s predicted probability distribution based on a watermark key $k$ to obtain a new probability distribution. Formally, given an LLM $P_{M}$ and a key $k$, the watermark rule $F$ modifies the distribution as follows:
$P_{M}^{F}(y_{i}|x,y_{1:i-1},k)=F(P_{M}(y_{i}|x,y_{1:i-1}),k)$ (1)
where $P_{M}^{F}$ is the modified probability distribution for the next token $y_{i}$.
The main difference between watermarking algorithms lies in how they determine the watermark key. Based on this, we categorize watermarking algorithms into n-gram based watermarking and fixed-key-list based watermarking. We will now introduce these two types of watermarking algorithms.
Definition 3 (N-Gram Based Watermarking).
In n-gram based watermarking, the watermark key $k_{i}$ for generating the current token $y_{i}$ is determined by a function $f$ that takes two inputs:
$k_{i}=f(K,y_{i-n:i-1})$ (2)
where $K$ is a pre-selected master key, and $y_{i-n:i-1}$ represents the previous $n$ tokens.
N-gram based watermarking ensures that for the same n-token prefix, the watermark key for generating the next token remains consistent. This approach is widely used in current watermarking algorithms, including KGW (Kirchenbauer et al., 2023a), KGW-V2 (Kirchenbauer et al., 2023b), Aar (Aaronson & Kirchner, 2022), DiPmark (Wu et al., 2023b), and SIR (Liu et al., 2024b). Next, we define fixed-key-list-based watermarking:
Definition 4 (Fixed-Key-List Based Watermarking).
Let $K=\{k_{1},k_{2},...,k_{m}\}$ be a fixed key list. For a given starting index $s\in\{1,...,m\}$, the watermark key $k_{i}$ for generating the $i$-th token is:
$k_{i}=k_{((s+i-1)\bmod m)+1}$ (3)
where the starting index $s$ may be randomly chosen for each generation process.
This approach is employed in algorithms such as Exp-Edit (Kuditipudi et al., 2023), where keys are used sequentially from a potentially random starting position in the key list.
To formalize our approach, we define our goal as follows:
Definition 5 (Black-box Watermark Identification).
A function $\mathcal{D}:P_{M}\rightarrow\{0,1\}$ that classifies a language model $P_{M}$ as watermarked (1) or not (0), without access to its internal parameters.
4 Watermarked LLM Identification
In this section, we introduce our Water-Probe algorithm, which identifies watermarks in LLMs through well-designed prompts. We first analyze why watermarked LLM identification is possible, then present the general algorithm process as shown in Fig. 1, and finally describe two specific algorithm variants.
4.1 Why Watermarked LLM Identification is Possible
Definition 2 implies that LLMs typically introduce some distortion to the distribution. However, there exist distortion-free watermarking algorithms that satisfy Definition 2. We first provide a definition and then demonstrate that LLMs with distortion-free watermarks can still be identified.
Definition 6 (Distortion-Free Watermark).
A watermarking algorithm is considered distortion-free if, for all possible inputs $x$ and partial output sequences $y_{1:i-1}$, the expected output distribution of the watermarked model $P_{M}^{F}$ over all possible watermark keys $k\in\mathcal{K}$ is identical to the original model $P_{M}$:
$\mathbb{E}_{k\in\mathcal{K}}[P_{M}^{F}(y_{i}|x,y_{1:i-1},k)]=P_{M}(y_{i}|x,y_{1:i-1}).$ (4)
This equation indicates that the expected output distribution of watermarked text remains unchanged when the watermark key is randomly selected across all possible keys. However, as shown in Equation 1, sampling with a specific watermark key $k$ introduces a difference between $P_{M}^{F}(y_{i}|x,y_{1:i-1},k)$ and $P_{M}(y_{i}|x,y_{1:i-1})$. This observation leads to the following theorem on detectability of watermarked LLMs:
Observation 1 (Distributional Difference of Watermarked LLMs).
Let $P_{M}$ be a language model and $F$ a watermark rule as defined in Definition 2. For a given watermark key $k$, the probability distribution of the watermarked model $P_{M}^{F}(y_{i}|x,y_{1:i-1},k)$ differs from the original distribution $P_{M}(y_{i}|x,y_{1:i-1})$. This distributional difference suggests the potential detectability of the watermark.
4.2 Pipeline of Watermarked LLM Identification
Observation 1 implies that the key to identifying a watermarked LLM is to construct prompts that allow for multiple samplings using the same watermark key to reveal the difference between $P_{M}^{F}(y_{i}|x,y_{1:i-1},k)$ and $P_{M}(y_{i}|x,y_{1:i-1})$. However, due to the black-box setting, we cannot directly access the origin logits $P_{M}(y_{i}|x,y_{1:i-1})$. Instead, we calculate the difference in LLM outputs for two distinct keys, defined as $\Delta(p,k_{m},k_{n})=P_{M}^{F}(y_{1}|p,k_{m})-P_{M}^{F}(y_{1}|p,k_{n})$, to aid in identification.
If two prompts yield similar output distributions, the same watermark key should have similar effects on both prompts (proven in Theorem 1). We determine if an LLM contains a watermark by comparing the consistency of the effects of two watermark keys on two similar prompts. Specifically, given $p_{1}$, $p_{2}$, $k_{1}$, and $k_{2}$, we assess the similarity between $\Delta(p_{1},k_{1},k_{2})$ and $\Delta(p_{2},k_{1},k_{2})$. High similarity indicates the presence of a watermark; otherwise, we conclude there is no watermark. Based on the above analysis, we now present the process of the Water-Probe algorithm.
Step 1: Construct highly correlated prompts. Construct $N$ prompts $P_{1},P_{2},...,P_{N}$ such that their output probability distributions under $M$ are highly similar, which can be expressed as:
$\forall i,j\in\{1,2,...,N\},\text{KL}(P_{M}(\cdot|P_{i})||P_{M}(\cdot|P_{j}))\leq\epsilon\text{ and }P_{i}\neq P_{j}$ (5)
[…]
Step 2: Sampling with simulated fixed watermark keys. Since we cannot access the logits under a given watermark key, we need to use repeated sampling to estimate the distribution. We construct a set of simulated watermark keys $K=\{k_{1},k_{2},...,k_{m}\}$ based on our prompt design (detailed in subsequent sections). For each prompt $P_{i}$ and each simulated key $k_{j}\in K$, we estimate the distribution as follows:
$\hat{P}_{M}^{F}(y|P_{i},k_{j})=\frac{1}{W}\sum_{w=1}^{W}\mathbf{1}_{y_{i,j}^{w}=y},\quad\text{where }y_{i,j}^{w}\sim P_{M}^{F}(y|P_{i},k_{j})$ (6)
where $W$ is the sample count, $\mathbf{1}_{A}$ is the indicator function, and $y_{i,j}^{w}$ is the $w$-th sample sampled from $P_{M}^{F}(y|P_{i},k_{j})$. Specific prompt techniques for different watermarking algorithms will be detailed later. Note that our prompt design for simulating watermark keys assumes the target LLM has a watermark. If it doesn’t, then $P_{M}^{F}(y|P_{i},k_{j})=P_{M}(y|P_{i})$ for all simulated keys.
Step 3: Analyze Cross-Prompt Watermark Consistency. We first assume that the watermark rule satisfies Lipschitz continuity. Based on this assumption, we can deduce that the differences in output distributions produced by a watermark key pair for highly correlated prompts are similar.
Assumption 1 (Lipschitz Continuity of Watermark Rule).
For prompts $p_{1}$ and $p_{2}$ satisfying the similarity condition in Equation 5, the watermark rule $F$ satisfies Lipschitz continuity. That is, there exists a constant $L>0$ such that for the probability distributions $P_{M}(\cdot|p_{1}),P_{M}(\cdot|p_{2})$ and any watermark key $k\in\mathcal{K}$:
$\|F(P_{M}(\cdot|p_{1}),k)-F(P_{M}(\cdot|p_{2}),k)\|_{1}\leq L\cdot\|P_{M}(\cdot|p_{1})-P_{M}(\cdot|p_{2})\|_{1}.$ (7)
Theorem 1 (Consistency of Watermark Effect).
Let $p_{1}$ and $p_{2}$ be two different prompts satisfying the similarity condition in Equation 5. Let $k_{1}$ and $k_{2}$ be two randomly sampled watermark keys from the key space $\mathcal{K}$. The effect of applying these keys on the output distribution should be highly consistent across prompts:
$\mathbb{E}_{k_{1},k_{2}\sim\mathcal{K}}[\text{Sim}(P_{M}^{F}(\cdot|p_{1},k_{1})-P_{M}^{F}(\cdot|p_{1},k_{2}),P_{M}^{F}(\cdot|p_{2},k_{1})-P_{M}^{F}(\cdot|p_{2},k_{2}))]\geq\rho$ (8)
where $P_{M}^{F}$ is the watermarked distribution, and $\text{Sim}(\cdot,\cdot)$ is a similarity measure (e.g., cosine similarity), and $\rho$ should be a constant significantly greater than 0.
Theorem 1 implies that for prompts with similar output distributions under the LLM, the expected differences introduced by the two watermark keys should also be similar. The formal proof is provided in Appendix A.
Based on Theorem 1, we calculate the average similarity using the estimated distributions from Step 2. To ensure stability across different sampling temperatures, we first apply a rank transformation. Specifically, for a token $y$, its rank is defined as the number of tokens with probability greater than or equal to that of $y$, denoted as $R(P)(y|x)=|\{y^{\prime}\in V:P(y^{\prime}|x)\geq P(y|x)\}|$. We then compute the expected similarity:
$\bar{S}=\frac{1}{N}\sum_{P_{i}\neq P_{j}\in\mathcal{P}}\sum_{k_{m}\neq k_{n}\in\mathcal{K}}\text{Sim}(\Delta_{R}(P_{i},k_{m},k_{n}),\Delta_{R}(P_{j},k_{m},k_{n}))$ (9)
Here, $\mathcal{P}$ is the prompt set, $\mathcal{K}$ is the watermark key set, $N=|\mathcal{P}|(|\mathcal{P}|-1)|\mathcal{K}|(|\mathcal{K}|-1)$, and $\Delta_{R}(P,k_{m},k_{n})=R(\hat{P}_{M}^{F}(\cdot|P,k_{m}))-R(\hat{P}_{M}^{F}(\cdot|P,k_{n}))$. We verify the importance of rank transformation in Appendix G.
According to Theorem 1, if $M$ contains a watermark, the similarity obtained from Equation 9 should be significantly greater than 0. If $M$ does not contain a watermark, we assume $P_{M}^{F}=P_{M}$ for any $k$, so Equation 9 should represent the similarity between two random vectors with zero mean, which should be close to 0. A detailed analysis of the no-watermark case is provided in Appendix B.
Based on this, we design the following z-test to perform hypothesis testing on the average similarity:
$z=(\bar{S}-\mu)/\sigma$ (10)
where $\bar{S}$ is the observed average similarity, $\sigma$ is the standard deviation of the $\bar{S}$, and $\mu$ is the mean of the $\bar{S}$ under the no-watermark case. Theoretically, $\mu$ should be 0 for an unwatermarked LLM. However, in practice, we may choose a value slightly greater than 0 to account for potential biases introduced by our prompt construction method or other factors. The standard deviation $\sigma$ is estimated through multiple experiments.
We reject the null hypothesis (no watermark) and conclude the LLM is likely watermarked if: $z>z_{\alpha}$, where $z_{\alpha}$ is the chosen significance level $\alpha$. In this work, we consider a z-score between 4 and 10 as moderate confidence, and above 10 as high confidence.
4.3 Constructing Repeated Sampling With Same Watermark Key
In the previous section, we introduced the basic pipeline of our Water-Probe algorithm. A key challenge is constructing prompts that enable multiple samplings with the same watermark key. This varies for different watermarking algorithms. We’ll discuss approaches for n-gram based and fixed-key-list based methods.
For N-gram based watermarking (Definition 3), since the watermark key is derived from the previous N tokens, we can design prompts that make the LLM generate N irrelevant tokens before following the prompt. An example is provided below:
  • Prompt 1: Example Prompt for Watermark-Probe-v1 Please generate abcd before answering the question. Question: Name a country with a large population. Answer: abcd India
In the example above, we assume generating abcd does not affect the distribution for answering the question. However, in practice, it’s challenging to ensure completely irrelevant tokens. Consequently, the $\bar{S}$ for an unwatermarked LLM constructed this way may be slightly above 0. We refer to the Watermark-Probe using prompts similar to the above table as Watermark-Probe-v1.
For fixed-key-set based watermarking, since the start watermark key is randomly selected each time, our approach is to approximate multiple samplings with the same watermark key by exploiting the correlation between the watermark key and the generated tokens. Specifically, we prompt the LLM to perform some quasi-random generation initially. Generally, the same watermark start key will only generate a few fixed sampling results, so we can assume that identical sampling result prefixes are generated by the same watermark key. Here’s a specific example of how we construct prompts for this approach:
  • Prompt 2: Example Prompt for Watermark-Probe-v2 Please generate a sentence that satisfies the following conditions: The first word is randomly sampled from A-Z. The second word is randomly sampled from zero to nine. The third word is randomly sampled from cat, dog, tiger and lion. Then answer the question: Name a country with a large population. Answer: A one cat China
As shown in the example above, we first prompt the LLM to generate N (3 in this case) random tokens before answering the question. Different random token combinations typically correspond to a few watermark keys. Figure 2 illustrates the distribution of watermark key counts for varying numbers of random tokens. As evident from the figure, given a fixed prefix, the vast majority of cases utilize a specific key. So this approach thus approximates sampling with the same watermark key. Similarly, we refer to the prompting method in the above table as Watermark-Probe-v2.
We provide the detailed steps of the Water-Probe algorithm in Algorithm 1 in the appendix.
Figure 2: Distribution of start keys for identical prefixes in KTH watermarking. Analysis based on prompts described in Section 4.3 for Watermark-Probe-v2. Each subplot represents a specific prefix(in title).
5 Experiment On Watermarked LLM Identification
5.1 Experiment Setup
Table 1: Detection similarities for various LLMs with and without different watermarks, calculated using Equation 9 and our two identification methods: Water-Probe-v1 and Water-Probe-v2. indicates high-confidence watermark identification and indicates low-confidence watermark identification while no color indicates no watermark identified. The corresponding z-scores can be found in Table 7 in the Appendix.
LLM N-Gram Fixed-Key-List
Non KGW Aar KGW-Min KGW-Skip DiPmark $\gamma$-Reweight EXP-Edit ITS-Edit
Water-Probe-v1 (w. prompt 4.3)
Qwen2.5-1.5B 0.02 ± 0.02 0.37 ± 0.02 0.88 ± 0.06 0.37 ± 0.02 0.39 ± 0.01 0.55 ± 0.01 0.55± 0.01 0.01± 0.02 0.00± 0.04
OPT-2.7B 0.05 ± 0.01 0.47 ± 0.01 0.91 ± 0.01 0.42 ± 0.02 0.45 ± 0.01 0.60 ± 0.01 0.61± 0.01 0.08 ± 0.02 0.09± 0.01
Llama-3.2-3B 0.04 ± 0.02 0.53 ± 0.01 0.90 ± 0.01 0.48 ± 0.00 0.49 ± 0.01 0.61 ± 0.01 0.61± 0.01 0.03 ± 0.01 0.04± 0.01
Qwen2.5-3B 0.03 ± 0.01 0.33± 0.02 0.75± 0.05 0.33± 0.02 0.38 ± 0.00 0.51± 0.01 0.53± 0.01 0.03± 0.01 0.06± 0.02
Llama2-7B 0.02 ± 0.01 0.42 ± 0.01 0.87 ± 0.01 0.31 ± 0.01 0.42 ± 0.01 0.56 ± 0.01 0.56± 0.04 0.03 ± 0.02 0.02± 0.00
Mixtral-7B 0.01 ± 0.02 0.41 ± 0.01 0.85 ± 0.02 0.37 ± 0.01 0.41 ± 0.02 0.57 ± 0.01 0.58± 0.03 0.00± 0.00 0.02± 0.02
Qwen2.5-7B 0.07 ± 0.04 0.41 ± 0.02 0.82 ± 0.02 0.34 ± 0.03 0.38 ± 0.02 0.43 ± 0.03 0.43± 0.02 0.06± 0.01 0.04± 0.02
Llama-3.1-8B 0.01 ± 0.02 0.41 ± 0.02 0.85 ± 0.02 0.41 ± 0.01 0.39 ± 0.01 0.57 ± 0.02 0.58± 0.00 0.02± 0.02 0.00± 0.01
Llama2-13B 0.01 ± 0.03 0.41 ± 0.01 0.86 ± 0.01 0.31 ± 0.02 0.40 ± 0.02 0.58 ± 0.02 0.60± 0.01 0.02 ± 0.01 0.02± 0.03
Average 0.029 0.418 0.854 0.371 0.412 0.553 0.505 0.031 0.032
Water-Probe-v2 (w. prompt 4.3)
Qwen2.5-1.5B 0.02± 0.02 0.30± 0.01 0.83± 0.01 0.29± 0.01 0.27± 0.02 0.49± 0.02 0.52± 0.03 0.39± 0.03 0.60 ± 0.00
OPT-2.7B 0.04 ± 0.03 0.29 ± 0.02 0.88 ± 0.01 0.23 ± 0.01 0.19 ± 0.02 0.42 ± 0.01 0.43± 0.03 0.43 ± 0.01 0.62 ± 0.00
Llama-3.2-3B 0.00 ± 0.01 0.31 ± 0.01 0.89 ± 0.01 0.33 ± 0.00 0.24 ± 0.01 0.51 ± 0.01 0.54± 0.01 0.52 ± 0.01 0.84 ± 0.00
Qwen2.5-3B 0.03± 0.02 0.35± 0.04 0.78± 0.01 0.29± 0.02 0.28± 0.01 0.45± 0.02 0.45± 0.02 0.39± 0.02 0.71 ± 0.00
Llama2-7B 0.04 ± 0.02 0.34 ± 0.01 0.82 ± 0.02 0.33 ± 0.01 0.28 ± 0.01 0.50 ± 0.01 0.51± 0.02 0.48 ± 0.01 0.81 ± 0.00
Mixtral-7B 0.09 ± 0.01 0.34 ± 0.04 0.83 ± 0.01 0.29 ± 0.02 0.24 ± 0.01 0.51 ± 0.01 0.53± 0.00 0.42± 0.02 0.81 ± 0.00
Qwen2.5-7B -0.01 ± 0.04 0.26 ± 0.02 0.70 ± 0.00 0.28 ± 0.02 0.23 ± 0.01 0.32 ± 0.03 0.35± 0.02 0.32± 0.02 0.73 ± 0.00
Llama-3.1-8B 0.01 ± 0.00 0.31 ± 0.01 0.77 ± 0.01 0.29 ± 0.02 0.26 ± 0.00 0.50 ± 0.01 0.51 ± 0.01 0.43± 0.01 0.71 ± 0.00
Llama2-13B 0.01 ± 0.02 0.35 ± 0.01 0.82 ± 0.02 0.26 ± 0.02 0.26 ± 0.01 0.50 ± 0.01 0.53± 0.01 0.44 ± 0.02 0.73 ± 0.00
Average 0.026 0.317 0.813 0.288 0.250 0.467 0.486 0.424 0.729
Tested Watermarking Algorithms: We evaluated a diverse range of LLM watermarking algorithms, including N-Gram based watermarking and Fixed-Key-List based watermarking. For N-Gram based watermarking, we tested KGW (Kirchenbauer et al., 2023a) ($\gamma=0.5$, $\delta=2$), Aar (Aaronson & Kirchner, 2022) ($N=1$), KGW-Min Kirchenbauer et al. (2023b) (window size of 4), KGW-Skip (Kirchenbauer et al., 2023b) (window of 3), DiPMark (Wu et al., 2023b) ($\alpha=0.45$), and $\gamma$ reweighting (Hu et al., 2023). For Fixed-Key-List based watermarking, we examined EXP-Edit (Kuditipudi et al., 2023) and ITS-Edit (Kuditipudi et al., 2023), both with a key length of 420. Details of these algorithms are provided in Appendix F.
Tested LLMs: To comprehensively evaluate our algorithm’s effectiveness, we tested a diverse range of LLMs with varying parameter sizes, including Qwen2.5-1.5B (Hui et al., 2024), OPT-2.7B (Zhang et al., 2022), Llama3.2-3B (Meta AI, 2024), Qwen2.5-3B, Llama2-7B (Touvron et al., 2023), Mixtral-7B (Jiang et al., 2024a), Qwen2.5-7B, Llama-3.1-8B (Dubey et al., 2024), and Llama2-13B (Touvron et al., 2023).We evaluated our Water-Probe algorithm on all LLMs, testing its performance under various watermarking schemes and in scenarios without watermarks.
Watermark-Probe Settings: For our Watermark-Probe algorithm, detailed prompts are provided in the appendix C. To calculate the z-score, we repeat each detection experiment 3 times to compute the standard deviation. We set $\mu=0.1$ for our experiments.
5.2 Main Results
In Table 1, we present the average similarity and standard deviation obtained using Watermark-Probe-v1 and Watermark-Probe-v2 algorithms for identifying various LLMs under different watermarking conditions and non-watermarked scenarios. For all LLMs, the sampling temperature was set to 1, with the number of samples set to $10^{4}$. As evident from Table 1, Watermark-Probe-v1 demonstrates high effectiveness for N-gram based watermarking but is not applicable to fixed-key-list based watermarking. In contrast, the Watermark-Probe-v2 algorithm proves effective in identifying all watermarking algorithms tested. Additionally, even watermarking algorithms claiming to be distortion-free, such as Aar (Aaronson & Kirchner, 2022), DiPMark (Wu et al., 2023b), and $\gamma$-reweighting (Hu et al., 2023), they can be effectively identified by both versions of Watermark-Probe. Furthermore, our algorithm maintains low similarity for non-watermarked LLMs, ensuring minimal false positive rates. Additionally, we calculated the average similarity for different watermarking algorithms in Table 1 to demonstrate their detection confidence. Among these, the Aar algorithm is the most easily detectable due to its pronounced perturbation for individual keys. Lastly, given the same number of samples, the Watermark-Probe-v1 algorithm yields more significant identification results for N-gram based watermarking algorithms compared to the Watermark-Probe-v2 algorithm.
Figure 3: The left plot shows the variation of z-scores detected by Watermark-Probe-v1 and Watermark-Probe-v2 as a function of sampling temperature. The right plot illustrates the change in z-scores detected by Watermark-Probe-v1 and Watermark-Probe-v2 with different sampling numbers.
5.3 Further Analysis
Influence of Sampling Temperature: The results in Table 1 are based on a sampling temperature of 1. To further validate the performance of Watermark-Probe under different sampling temperatures, we show the z-score changes across temperatures in the left plot of Figure 3. Using Llama2-7B as an example, as the sampling temperature increases from 0.1 to 1.5, both Watermark-Probe-v1 and Watermark-Probe-v2 can distinguish between watermarked and unwatermarked LLMs. However, at relatively low temperatures ($T<0.5$), detection of unwatermarked LLMs may show some fluctuations. Since deployed LLMs rarely use very low temperatures, our algorithm can be considered effective for detecting real-world LLM deployments.
Influence of Sampling Number: The right plot in Figure 3 illustrates the impact of the sampling number on the detected z-score. Specifically, we used Llama2-7B as the target LLM with a temperature of 1. We observed that Watermark-Probe-v2 requires more samples compared to Watermark-Probe-v1. With insufficient samples, Watermark-Probe-v2 lacks enough common prefixes to compute Equation 9. In our setting, Watermark-Probe-v1 can achieve stable detection with 1,000 samples, while Watermark-Probe-v2 requires at least $10^{4}$ samples. For cases where detection is successful, the z-score of watermarked LLMs tends to increase with the number of samples, although this trend exhibits fluctuations.
6 Enhancing the Imperceptibility of Watermarked LLMs
We have demonstrated that current watermarked LLMs can be identified by our Water-Probe method. In this section, we discuss how to improve the imperceptibility of watermarked LLMs. The core principle is to make it challenging to construct repeated sampling scenarios using two separate keys according to Equation 9.
One design is globally fixed watermark key (e.g., Unigram (Zhao et al., 2023)). While it’s easy to construct repeated sampling scenarios with a single key, we cannot detect stable deviations between different keys as only one exists globally. In the appendix E, we provide an algorithm named Water-Contrast to identify watermarks by comparing the target LLM distribution and a prior distribution. While not theoretically guaranteed (it’s challenging to determine if this bias is from watermarking or inherent to the LLM), it shows practical effectiveness. Meanwhile, Unigram watermarks are susceptible to cracking (Jovanović et al., 2024).
The second design aims to increase the randomness of watermark key selection, making it less dependent on N-grams. This approach makes it difficult to construct repeated sampling scenarios using the same key. For Fixed-Key-List Based watermarking, a viable strategy is to increase the length of the key list. Since the initial key position is random, increasing the key list length enhances the randomness of key selection.
Figure 4: Distribution of start keys for identical prefixes in Water-Bag strategy. Analysis based on prompts described in Section 4.3 for Water-Probe-v2. Each subplot represents a specific prefix (showed in title).
Additionally, for N-gram Based watermarking algorithms, we propose an enhanced strategy called Water-Bag, which combines multiple master watermark keys into a key-bag with a key inversion mechanism.
Definition 7 (Water-Bag Strategy).
The Water-Bag strategy extends N-gram based watermarking by using a set of master keys $K=\{K_{1},K_{2},...,K_{n}\}$ and their inversions $\overline{K}=\{\overline{K_{1}},\overline{K_{2}},...,\overline{K_{n}}\}$. For each generation, a master key $K_{j}$ or its inversion $\overline{K_{j}}$ is randomly selected:
$P_{M}^{WB}(y_{i}|x,y_{1:i-1},K,\overline{K})=F(P_{M}(y_{i}|x,y_{1:i-1}),k_{i}),\quad k_{i}=f(K_{j}^{*},y_{i-n:i-1}),\quad K_{j}^{*}\sim\text{Uniform}(K\cup\overline{K})$
where $P_{M}^{WB}$ is the modified probability distribution, $K_{j}^{*}$ is randomly sampled from the combined set of original and inverted keys, and $f$ is the watermark key derivation function. The inverted key $\overline{K_{j}}$ is defined as:
$\frac{1}{2}(F(P_{M}(y_{i}|x,y_{1:i-1}),f(K_{j},y_{i-n:i-1}))+F(P_{M}(y_{i}|x,y_{1:i-1}),f(\overline{K_{j}},y_{i-n:i-1})))=P_{M}(y_{i}|x,y_{1:i-1})$
This ensures that the average effect of $K_{j}$ and $\overline{K_{j}}$ on the logits is equivalent to the original logits, which makes our Water-Probe-v1 ineffective against the Water-Bag strategy.
For the watermarked text detection for Water-Bag Strategy, we use the maximum detection score across all master keys in the bag. The text is considered watermarked if this maximum exceeds a threshold.
Table 2: Performance comparison of Water-Bag Strategy and Exp-Edit algorithm in watermarked LLM identification and watermarked text detection. Water-Bag is evaluated with varying bag sizes, while Exp-Edit is tested with different key lengths. represents high-confidence watermark and represents low-confidence watermark. Detail of z-score could be seen in Appendix H.
KGW w. Water-Bag Exp-Edit(Key-len)
None $|K\cup\overline{K}|=1$ $|K\cup\overline{K}|=2$ $|K\cup\overline{K}|=4$ $|K\cup\overline{K}|=8$ $|K|=420$ $|K|=1024$ $|K|=2048$
Watermarked LLM Indentification
Water-Probe-v1(n=3) 0.02 ± 0.01 0.42 ± 0.01 0.05 ± 0.01 0.02 ± 0.01 0.03 ± 0.02 0.03 ± 0.05 0.02 ± 0.01 0.02 ± 0.02
Water-Probe-v2(n=3) 0.04 ± 0.01 0.34 ± 0.01 0.34 ± 0.01 0.25 ± 0.01 0.16 ± 0.02 0.48 ± 0.01 0.33±0.01 0.23±0.02
Water-Probe-v2(n=5) 0.06 ± 0.06 0.32 ± 0.01 0.18 ± 0.01 0.12 ± 0.02 0.07 ± 0.01 0.64 ± 0.00 0.54 ± 0.01 0.44 ± 0.00
Watermarked Text Detection
Detection-F1-score - 1.0 1.0 1.0 1.0 1.0 0.975 1.0
PPL 8.15 11.93 11.85 12.17 12.50 16.63 17.28 19.06
Robustness (GPT3.5) - 0.843 0.849 0.748 0.696 0.848 0.854 0.745
Detection-time (s) - 0.045 0.078 0.156 0.31 37.87 108.5 194.21
To validate the effectiveness of the two strategies for enhancing the imperceptibility of watermarked LLMs, we evaluated their performance in Table 2. We examined both watermarked LLM identification and watermarked text detection settings, assessing the new watermarking strategies’ detectability by our Water-Probe algorithm and their impact on watermarked text detection efficacy and performance.
For watermarked text detection, we used OPT-2.7B to generate texts on the C4 dataset (Raffel et al., 2020), using 30 tokens as prompts and generating 200 additional tokens with watermarks. PPL was calculated using Llama2-7B. To assess detection robustness, we computed the F1-score after rewriting texts using GPT-3.5. Detection time for single text was also recorded. We use the KGW algorithm as an implementation example for the water-bag strategy. For watermarked LLM identification, we report results for both Water-Probe-v1 and Water-Probe-v2. For Water-Probe-v2, we present results for $n=3$ and $n=5$, where $n$ is the number of random tokens generated as described in Section 4.3. The $n=3$ setting matches Table 1, while prompts for $n=5$ are provided in Appendix C.
Table 2 demonstrates that Water-Probe-v1 fails to effectively identify both Water-Bag and Exp-Edit algorithms. For Water-Probe-v2, detection difficulty increases with larger bag sizes in Water-Bag and longer key lengths in Exp-Edit. However, Water-Bag proves more challenging to identify. Crucially, Water-Bag’s detectability remains stable as $n$ increases, while Exp-Edit becomes more easy to identify. We analyzed the key distribution of Water-Bag under Water-Probe-v2 for different prefixes in Figure 4. The distribution is notably more uniform compared to Exp-Edit in Figure 2, explaining why Water-Bag is relatively harder to identify. We observe that increasing watermark key randomness reduces watermark robustness for both strategies, with Exp-Edit also significantly increasing detection time. This highlights a trade-off in watermarking algorithms between key randomness and robustness (Liu et al., 2024c). Future work should focus on developing algorithms that enhance randomness without compromising robustness.
7 Conclusion
In this paper, we pioneered the study of identifying watermarked LLMs. We first theoretically demonstrated the basis for identifying watermarked LLMs. We then designed the Water-Probe algorithm, which identifies watermarked LLMs by comparing distribution differences of similar prompts under different watermark keys. Our experiments showed that our algorithm is applicable to all N-Gram and Fixed-Key-List based Watermarking algorithms, independent of sampling temperature. We discussed scenarios where Water-Probe might fail and designed the WaterBag watermarking algorithm, which sacrifices some robustness of watermarked text detection to make watermarked LLMs harder to identify. Future work could focus on watermark concealment as a key research direction, designing more covert watermarking schemes.
Appendix A Proof of Theorem 1 and Additional Analyses
Proof.
Goal: Prove that under the Lipschitz continuity condition, the watermark modification differences for two similar prompts $p_{1}$ and $p_{2}$ have high similarity, with an expected similarity of at least $\rho$.
Step 1: Express the watermark modification difference
For a prompt $p$ and keys $k_{1},k_{2}$, define the watermark modification difference as:
$\Delta_{p}(k_{1},k_{2})=P_{M}^{F}(\cdot|p,k_{1})-P_{M}^{F}(\cdot|p,k_{2})$ (13)
Step 2: Utilize the Lipschitz continuity condition
By the Lipschitz continuity condition, for any fixed key $k$:
$\|P_{M}^{F}(\cdot|p_{1},k)-P_{M}^{F}(\cdot|p_{2},k)\|_{1}\leq L\cdot\|P_{M}(\cdot|p_{1})-P_{M}(\cdot|p_{2})\|_{1}$ (14)
Since $p_{1}$ and $p_{2}$ satisfy the similarity condition $\text{KL}(P_{M}(\cdot|p_{1})||P_{M}(\cdot|p_{2}))\leq\epsilon$ (Equation 5), by Pinsker’s inequality (Csiszár & Körner, 2011):
$\|P_{M}(\cdot|p_{1})-P_{M}(\cdot|p_{2})\|_{1}\leq\sqrt{2\epsilon}$ (15)
Therefore, for any $k\in\{k_{1},k_{2}\}$:
$\|P_{M}^{F}(\cdot|p_{1},k)-P_{M}^{F}(\cdot|p_{2},k)\|_{1}\leq L\cdot\sqrt{2\epsilon}$ (16)
Step 3: Analyze the difference of watermark modification differences
Consider the definition of $\Delta_{p}(k_{1},k_{2})$ and compute $\|\Delta_{p_{1}}(k_{1},k_{2})-\Delta_{p_{2}}(k_{1},k_{2})\|_{1}$:
$\|\Delta_{p_{1}}(k_{1},k_{2})-\Delta_{p_{2}}(k_{1},k_{2})\|_{1}$ $=\|(P_{M}^{F}(\cdot|p_{1},k_{1})-P_{M}^{F}(\cdot|p_{1},k_{2}))-(P_{M}^{F}(\cdot|p_{2},k_{1})-P_{M}^{F}(\cdot|p_{2},k_{2}))\|_{1}$
$=\|(P_{M}^{F}(\cdot|p_{1},k_{1})-P_{M}^{F}(\cdot|p_{2},k_{1}))-(P_{M}^{F}(\cdot|p_{1},k_{2})-P_{M}^{F}(\cdot|p_{2},k_{2}))\|_{1}$
$\leq\|P_{M}^{F}(\cdot|p_{1},k_{1})-P_{M}^{F}(\cdot|p_{2},k_{1})\|_{1}+\|P_{M}^{F}(\cdot|p_{1},k_{2})-P_{M}^{F}(\cdot|p_{2},k_{2})\|_{1}$
$\leq L\cdot\sqrt{2\epsilon}+L\cdot\sqrt{2\epsilon}$
$=2L\cdot\sqrt{2\epsilon}$
$=\delta^{\prime}$ (17)
where $\delta^{\prime}=2L\cdot\sqrt{2\epsilon}$.
Step 4: Relate $L_{1}$ distance to similarity measure
Assume $\text{Sim}(\cdot,\cdot)$ is cosine similarity, which is negatively correlated with $L_{1}$ distance. Since $\|\Delta_{p_{1}}(k_{1},k_{2})-\Delta_{p_{2}}(k_{1},k_{2})\|_{1}\leq\delta^{\prime}$, and $\delta^{\prime}$ is a small positive number:
$\text{Sim}(\Delta_{p_{1}}(k_{1},k_{2}),\Delta_{p_{2}}(k_{1},k_{2}))\geq\rho^{\prime}$ (18)
where $\rho^{\prime}$ is a lower bound dependent on $\delta^{\prime}$, approaching 1 as $\delta^{\prime}$ decreases.
Step 5: Calculate expected similarity
Since $\text{Sim}(\Delta_{p_{1}}(k_{1},k_{2}),\Delta_{p_{2}}(k_{1},k_{2}))\geq\rho^{\prime}$ for any $k_{1},k_{2}\in\mathcal{K}$, for randomly sampled $k_{1},k_{2}$:
$\mathbb{E}_{k_{1},k_{2}\sim\mathcal{K}}\left[\text{Sim}(\Delta_{p_{1}}(k_{1},k_{2}),\Delta_{p_{2}}(k_{1},k_{2}))\right]\geq\rho^{\prime}$ (19)
Set $\rho=\rho^{\prime}$, and by choosing a sufficiently small $\epsilon$ (making $\delta^{\prime}$ small enough), we can ensure $\rho$ is close to 1.
Conclusion: Under the Lipschitz continuity condition, for two similar prompts $p_{1}$ and $p_{2}$, the expected similarity of their watermark modification differences under randomly sampled watermark keys $k_{1}$ and $k_{2}$ is at least $\rho$, where $\rho$ is a large positive number close to 1. This proves the high consistency of watermark effects in similar contexts, ensuring the detectability and robustness of the watermark. ∎
Appendix B Statistical Analysis of Unwatermarked LLMs Identification
When there’s no watermark present in the LLM, we expect the similarity measure in Equation 9 to be close to 0. This can be mathematically explained as follows:
For an unwatermarked model, $P_{M}^{F}=P_{M}$ for any $k$. Therefore:
$\Delta_{R}(P_{i},k_{m},k_{n})$ $=R(\hat{P}_{M}^{F}(\cdot|P_{i},k_{m}))-R(\hat{P}_{M}^{F}(\cdot|P_{i},k_{n}))$ (20)
$=R(\hat{P}_{M}(\cdot|P_{i}))-R(\hat{P}_{M}(\cdot|P_{i}))$
$=0$
Consequently, $\text{Sim}(\Delta_{R}(P_{i},k_{m},k_{n}),\Delta_{R}(P_{j},k_{m},k_{n}))\approx\text{Sim}(0,0)=0$ for any $P_{i},P_{j},k_{m},k_{n}$. This results in the average similarity $\bar{S}$ in Equation 9 being approximately 0.
Appendix C Detailed Prompts for Simulating Watermark Keys
In this section, we provide details of the prompt used for repeated sampling with the same key, as introduced in Section 4.3.
For the Water-Probe-v1 algorithm, we use the following prompt pair as shown in Prompt C and Prompt C.
  • Prompt3: First prompt for Fixed Key List Based Watermarking(Water-Probe-v1) Please generate some text based on the following instructions(no other words): First generate the prefix D seven tiger. Then answer the question: Name a country with a large population. Example1: D seven tiger China Example2: D seven tiger India Example3:
  • Prompt4: First prompt for Fixed Key List Based Watermarking(Water-Probe-v1) Please generate some text based on the following instructions(no other words): First generate the prefix D seven tiger. Then answer the question: Name a country with a large area. Example1: D seven tiger China Example2: D seven tiger India Example3:
Here, we use a fixed prefix D seven tiger as an example. In the actual experiment, we used 50 different prefixes to simulate 50 potentially different watermark keys. For the experiment in Table 1, we performed 10,000 samplings, with each prefix sampled 200 times. The complete list of 50 prefixes is shown below.
All the prefix for Water-Probe-v1 Y three lion, G three lion, U six lion, A eight tiger, K four cat, N seven tiger, K three cat, H five dog, E zero lion, V three dog, W five dog, K one tiger, B two tiger, E two lion, U six dog, A two tiger, D two tiger, I nine dog, F three lion, C three dog, N five cat, L two dog, K zero tiger, E five dog, B five cat, X four tiger, U three dog, K nine dog, P one dog, H zero dog, V eight tiger, S three tiger, P seven cat, S six dog, Y nine cat, J one tiger, C five tiger, A zero lion, L eight dog, X eight dog, I two dog, C eight tiger, O three tiger, L one cat, M five tiger, P five cat, F seven cat, I zero cat, P two lion, L four cat
  • Prompt5: First prompt for Fixed Key List Based Watermarking(Water-Probe-v2) Please generate some text based on the following instructions(no other words): The first word is randomly sampled from A-Z. The second word is randomly sampled from zero to nine. The third word is randomly sampled from cat, dog, tiger and lion. Then add a separator | and answer the following question: Name a country with a large population. Example1: D seven tiger | United States Example2: J five dog | India Example3: R six cat | China Example4: T one tiger | Indonesia Example5: P seven cat | Pakistan Example6: G six cat | Russia Example7: R five tiger | India Example8: L nine cat | Mexico Example9: T four tiger | United States Example10: H three dog | Japan Example11: B three lion | Germany Example12:
  • Prompt6: Second prompt for Fixed Key List Based Watermarking(Water-Probe-v2) Please generate some text based on the following instructions(no other words): The first word is randomly sampled from A-Z. The second word is randomly sampled from zero to nine. The third word is randomly sampled from cat, dog, tiger and lion. Then add a separator | and answer the following question: Name a country with a large area. Example1: L three tiger | United States Example2: X three cat | India Example3: A six tiger | China Example4: W eight lion | Argentina Example5: D five dog | France Example6: P one cat | Russia Example7: E six tiger | Australia Example8: Z eight lion | Canada Example9: Q two tiger | United States Example10: A nine cat | Brazil Example11: V three dog | Russia Example12:
Similarly, for Prompt C and Prompt C, we assume that the answer spaces for the questions Name a country with a large population. and Name a country with a large area are highly correlated, i.e., countries with large areas tend to have relatively large populations. In practice, other correlated prompt pairs can be chosen, as long as they satisfy the correlation requirement.
For the Water-Probe-v2 algorithm, we use the following prompt pair shown in Prompt C and Prompt C.
  • Prompt7: First prompt for Fixed Key List Based Watermarking(Water-Probe-v2 N=5) Please generate some text based on the following instructions(no other words): The first word is randomly sampled from A-Z. The second word is randomly sampled from zero to nine. The third word is randomly sampled from cat, dog, tiger and lion. The fourth word is randomly sampled from apple, banana and orange. The fifth word is randomly sampled from car, bus and truck. The sixth entry is the answer to the following question: Name a country with a large population. Example1: D seven tiger apple car United States Example2: J five dog banana bus India Example3: R six cat orange truck China Example4: T one tiger apple bus Indonesia Example5: P seven cat orange car Pakistan Example6: G six cat banana truck Russia Example7: R five tiger apple bus India Example8: L nine cat banana car Mexico Example9: T four tiger orange truck United States Example10: H three dog apple bus Japan Example11: B three lion orange truck Germany Example12:
  • Prompt8: Second prompt for Fixed Key List Based Watermarking(Water-Probe-v2 N=5) Please generate some text based on the following instructions(no other words): The first word is randomly sampled from A-Z. The second word is randomly sampled from zero to nine. The third word is randomly sampled from cat, dog, tiger and lion. The fourth word is randomly sampled from apple, banana and orange. The fifth word is randomly sampled from car, bus and truck. The sixth entry is the answer to the following question: Name a country with a large population. Example1: D seven tiger apple car United States Example2: J five dog banana bus India Example3: R six cat orange truck China Example4: T one tiger apple bus Argentina Example5: P seven cat orange car France Example6: G six cat banana truck Russia Example7: R five tiger apple bus Australia Example8: L nine cat banana car Canada Example9: T four tiger orange truck United States Example10: H three dog apple bus Brazil Example11: B three lion orange truck Russia Example12:
In this case, we selected the same question pair as in Water-Probe-v1. However, for Water-Probe-v2, all prefixes are randomly sampled by the LLM during the generation process. This random sampling by the LLM itself makes it easier to model the correlations between watermark keys at different positions.
Prompt C and Prompt C both generate prefixes of length 3. Since we further analyzed the case of generating prefixes of length 5 in Section 6, we also provide prompts for prefixes of length 5 in Prompt C and Prompt C. It is worth noting that the longer the required prefix, the more total sampling times are needed, as more samples are required to cover all actually occurring prefixes. In this work, for the case where the prefix length is 3, we typically sampled 10,000 times, and for the case where the prefix length is 5, we typically sampled 100,000 times.
Appendix D Detailed Algorithm for Water-Probe
To provide a clearer presentation of our Water-Probe algorithm, we present here a complete algorithmic representation, corresponding to the algorithm pipeline process described in Section 4.2. This algorithm flow can be used for both Water-Probe-v1 and Water-Probe-v2, although the prompt construction process differs between them, as detailed in Appendix C.
Algorithm 1 Water-Probe Algorithm
1:LLM $M$, significance level $\alpha$, sampling count $W$
2:Watermark detection result
3:Construct highly correlated prompts $P_{1},P_{2},...,P_{N}$ satisfying Equation 5
4:Define watermark key list $K=\{k_{1},k_{2},...,k_{m}\}$
5:for each prompt $P_{i}$ and key $k_{j}\in K$ do
6:     Initialize count dictionary $C_{i,j}(y)\leftarrow 0$ for all $y\in V$
7:     for $w=1$ to $W$ do
8:         Construct sampling prompt $P_{i,j}^{w}$ based on watermarking method (see Section 4.3)
9:         Generate output $y_{i,j}^{w}\sim P_{M}^{F}(\cdot|P_{i,j}^{w},k_{j})$
10:         $C_{i,j}(y_{i,j}^{w})\leftarrow C_{i,j}(y_{i,j}^{w})+1$
11:     end for
12:     $\hat{P}_{M}^{F}(y|P_{i},k_{j})\leftarrow C_{i,j}(y)/W$ for all $y\in V$
13:end for
14:Apply rank transformation to all $\hat{P}_{M}^{F}(\cdot|P_{i},k_{j})$
15:Calculate average similarity $\bar{S}$ using Equation 9
16:Compute z-score: $z=(\bar{S}-\mu)/\sigma$
17:if $z>z_{\alpha}$ then
18:     return LLM is likely watermarked
19:else
20:     return No evidence of watermarking
21:end if
Appendix E Detection of Global-Fixed Key Watermarking in LLMs
In this section, we explore how to identify LLMs with global-fixed key based watermarking. As discussed in Section 6, since global-fixed key based watermarking uses only one key globally, we cannot compare differences between two different watermark keys. However, as the global-fixed key produces consistent bias across all prompts, we can calculate a prior distribution for each prompt and then verify if the differences from this prior distribution have high similarity across all prompt lists satisfying Equation 5.
Specifically, let $P_{1},P_{2},...,P_{N}$ be the constructed prompt list, and $P_{\text{prior}}$ be their prior distribution. We can modify the average similarity calculation in Equation 9 as follows:
Table 3: Identification of global fixed key watermarking (e.g., Unigram watermarking) in LLMs using a prior distribution computed as the average output distribution of all eight LLMs. indicates high-confidence watermark identification and indicates low-confidence watermark identification while no color indicates no watermark identified.
LLM Similarity Z-Score
Unigram Unwatermark Unigram Unwatermark
GPT2 0.59 ± 0.018 0.06 ± 0.37 32.49 1.61
OPT1.3B 0.6 ± 0.02 0.06 ± 0.029 28.77 2.07
OPT2.7B 0.65 ± 0.009 0.03 ± 0.048 74.81 0.63
LLama2-7B 0.48 ± 0.042 0.04 ± 0.043 11.31 0.92
LLama-2-13B 0.38 ± 0.037 0.08 ± 0.018 10.17 4.54
LLama-3.1-8B 0.68 ± 0.06 0.12 ± 0.05 12.09 2.42
Mixtral-7B 0.87 ± 0.056 0.038 ± 0.052 22.24 0.74
Qwen2.5-7B 0.53 ± 0.058 0.09 ± 0.028 8.96 3.29
Table 4: Identification of global fixed key watermarking in LLMs using a single proxy model’s output as prior distribution. indicates high-confidence watermark identification and indicates low-confidence watermark identification while no color indicates no watermark identified.
LLM GPT2 OPT1.3B OPT2.7B LLama2-7B
Similarity Z-Score Similarity Z-Score Similarity Z-Score Similarity Z-Score
W NW W NW W NW W NW W NW W NW W NW W NW
GPT2 0.48±0.009 0.005±0.025 49.68 0.21 0.58±0.004 0.07±0.02 138.24 3.46 0.65±0.008 -0.07±0.015 77.60 -0.49 0.42±0.017 0.034±0.015 25.37 2.25
OPT1.3B 0.71±0.07 0.09±0.04 9.433 2.31 0.59±0.007 0.04 ±0.02 81.44 1.84 0.67±0.007 0.05±0.015 92.46 3.43 0.45±0.008 0.11±0.04 53.21 2.55
OPT2.7B 0.7±0.03 0.1±0.05 23.89 1.70 0.59±0.01 0.07±0.03 53.05 2.41 0.64±0.007 0.025±0.025 88.79 0.99 0.46±0.04 0.03±0.01 11.16 2.47
LLama2-7B 0.66±0.015 0.1±0.007 44.03 14.08 0.57±0.025 0.07±0.02 22.38 3.32 0.67±0.008 0.028±0.023 79.72 1.21 0.43±0.02 0.06±0.02 18.35 2.60
  • Prompt9: Used Prompt Set for Global-Fixed key Based Watermarking 1 : Please generate random number sequence between 0 to 9:2,3,9,8,0,4,7,5,6,1,5,8,7,1,2,4,6,0,9,3,2 : Please generate random number sequence between 0 to 9:3,8,0,7,4,5,6,1,9,2,1,3,7,9,2,0,4,6,8,5,3 : Please generate random number sequence between 0 to 9:5,8,7,1,2,4,6,0,9,3,2,3,9,8,0,4,7,5,6,1,4 : Please generate random number sequence between 0 to 9:0,7,2,3,6,5,1,9,8,4,3,8,0,7,4,5,6,1,9,2,5 : Please generate random number sequence between 0 to 9:1,3,7,9,2,0,4,6,8,5,0,7,2,3,6,5,1,9,8,4,6 : Please generate random number sequence between 0 to 9:6,0,9,3,2,3,9,8,0,4,7,5,6,1,5,8,7,1,2,4,7 : Please generate random number sequence between 0 to 9:4,6,0,9,3,2,3,9,8,0,4,7,5,6,1,5,8,7,1,2,8 : Please generate random number sequence between 0 to 9:9,8,0,4,7,5,6,1,5,8,7,1,2,4,6,0,9,3,1,3,9 : Please generate random number sequence between 0 to 9:7,4,5,6,1,9,2,1,3,7,9,2,0,4,6,8,5,0,7,2,10 : Please generate random number sequence between 0 to 9:8,0,7,4,5,6,1,9,2,5,8,7,1,2,4,6,0,9,3,2,
$\bar{S}=\frac{1}{|\mathcal{P}|(|\mathcal{P}|-1)}\sum_{P_{i}\neq P_{j}\in\mathcal{P}}\text{Sim}(R(P_{M}(\cdot|P_{i})-P_{\text{prior}}(\cdot|P_{i})),R(P_{M}(\cdot|P_{j})-P_{\text{prior}}(\cdot|P_{j}))).$ (21)
For the prior distribution, we select N proxy LLMs and calculate the average output probability distribution under prompt $p_{i}$ as the prior distribution. Specifically:
$P_{\text{prior}}(\cdot|p_{i})=\frac{1}{N}\sum_{j=1}^{N}P_{M_{j}}(\cdot|p_{i}),$ (22)
where $M_{j}$ represents the $j$-th proxy LLM, and $N$ is the total number of proxy LLMs used. Specifically, we used the following proxy LLMs: GPT2, OPT1.3B, OPT2.7B, Llama2-7B, Llama2-13B, Llama3-8.1B, Mixtral-7B, and Qwen2.5-7B.
Additionally, we utilized 10 distinct prompts for computation to validate our method as shown in Prompt E.
Table 5: Details of watermarking algorithms tested in our work.
Algorithm Name Category Methodology
KGW (Kirchenbauer et al., 2023a) N-Gram Separate the vocabulary set into two lists: a red list and a green list based on the preceding token, then add bias to the logits of green tokens so that the watermarked text exhibits preference of using green tokens.
KGW-Min (Kirchenbauer et al., 2023b) N-Gram Similar to KGW, this approach partitions the vocabulary set based on the minimum token ID within a window of N-gram preceding tokens.
KGW-Skip (Kirchenbauer et al., 2023b) N-Gram Similar to KGW, this approach partitions the vocabulary set based on the left-most token ID within a window of N-gram preceding tokens.
Aar (Aaronson & Kirchner, 2022) N-Gram Generate a pseudo-random vector $r_{t}$ based on the N-gram preceding tokens to guide sampling at position $t$, and choose the token $i$ that maximize $r_{t}(i)^{1/p_{t}(i)}$ (exponential minimum sampling), where $p_{t}$ is the probability produced by LLM.
$\gamma$-Reweight (Hu et al., 2023) N-Gram Randomly shuffle the probability vector using a seed based on the preceding N-gram tokens. Discard the left half of the vector, doubling the remaining probabilities. Conduct further sampling using this reweighted distribution.
DiPmark (Wu et al., 2023b) N-Gram Similar to $\gamma$-Reweight, after shuffling, discard the left $\alpha$ portion of the vector and amplify the remaining probabilities by $1/(1-\alpha)$.
EXP-Edit (Kuditipudi et al., 2023) Fixed-Key-List Based on the Aar concept, construct a fixed pseudo-random vector list. When generating watermarked text, randomly select a start index in the list. For each watermarked token generation, sequentially use the pseudo-random vectors from this index for exponential minimum sampling. During detection, employ edit distance to calculate the correlation between the pseudo-random vector list and the text.
ITS-Edit (Kuditipudi et al., 2023) Fixed-Key-List Similar to EXP-Edit, this method also uses a fixed pseudo-random vector list. However, it uses inverse transform sampling instead of exponential minimum sampling during token selection.
The intuition behind this method is that if $M$ is watermarked, the differences between its output distributions and those of the proxy model should exhibit consistent patterns across different prompts, resulting in a high correlation. Conversely, for an unwatermarked model, we assume that the differences in bias between different language models are relatively small, and the correlation of these differences should be low.
Table 3 presents the identification results for global-fixed key watermarking using a prior distribution. Our method effectively identifies models employing global-fixed key watermarking, while yielding low z-scores for unwatermarked LLMs.
To further validate the key factors in using prior distribution for testing, we conducted experiments using a single proxy model as the prior distribution, as shown in Table 4. We performed cross-experiments with different LLMs. These results demonstrate that global fixed-key watermarking can still be detected when using a single LLM as the prior distribution. However, the z-score detection for unwatermarked LLMs exhibits greater fluctuation. This is primarily due to the significant bias in a single proxy model as the prior distribution, affecting the variance of identification and resulting in small z-score.
Although our method using prior distribution achieved good results in our experiments, this identification approach has a limitation in that it can only detect stable biases in LLMs, assuming that a stable bias indicates a watermark. However, this assumption may not hold in real-world scenarios, as it is challenging to distinguish whether the bias is caused by watermarking or inherent to the LLM itself. This is particularly problematic in cases where LLMs are known to have inherent biases. Future work could investigate more interpretable detection methods.
Table 6: Z-scores of waterbag method and Exp-Edit method. represents high-confidence watermark and represents low-confidence watermark, while no color means no watermark. This table provides supplementary information on the similarity content in Table 2.
KGW w. Water-Bag Exp-Edit(Key-len)
None $|K\cup\overline{K}|=1$ $|K\cup\overline{K}|=2$ $|K\cup\overline{K}|=4$ $|K\cup\overline{K}|=8$ 420 1024 2048
Watermarked LLM Indentification
Water-Probe-v1(n=3) -8.20 11.67 -4.22 -7.84 -4.34 -3.07 -5.87 -5.17
Water-Probe-v2(n=3) -2.87 24.87 23.49 16.19 3.08 47.74 18.70 6.17
Water-Probe-v2(n=5) -100.43 28.56 13.68 1.38 -4.41 131.30 78.63 69.40
Figure 5: The variation of z-scores at different temperatures when calculating similarity without using rank transformation in Equation 9.
Table 7: Z-scores for watermark detection on various LLMs with and without watermarks using Watermark-Probe-1 and Watermark-Probe-2. indicates high-confidence watermark identification, indicates low-confidence watermark identification, while no color indicates no watermark identified. This table provides supplementary information on the similarity content in Table 1.
LLM N-Gram Fixed-Key-Set
Non KGW Aar KGW-Min KGW-Skip DiPmark $\gamma$-Reweight EXP-Edit ITS-Edit
Watermark-Probe-1 (w. prompt 4.3)
Qwen2.5-1.5B -5.02 43.32 123.57 14.70 24.29 33.35 55.90 -5.21 -25.04
OPT-2.7B -5.99 49.21 117.95 16.69 35.78 93.25 49.98 -1.32 -1.27
Llama-3.2-3B -4.52 49.39 79.33 80.94 71.11 87.17 76.35 -6.70 -4.29
Qwen2.5-3B -6.70 48.20 127.07 14.73 625.36 56.15 42.96 -6.18 -1.97
Llama2-7B -8.20 30.01 109.00 25.51 29.75 44.35 106.37 -3.07 -28.05
Mixtral-7B -3.80 25.03 40.21 35.51 19.99 36.30 137.39 -22.01 -3.31
Qwen2.5-7B -1.16 38.25 30.03 22.85 50.34 47.31 50.48 -3.39 -15.92
Llama-3.1-8B -6.44 20.40 143.52 29.05 28.19 29.01 125.46 -3.79 -12.70
Llama2-13B -3.62 29.23 79.42 11.74 18.01 30.40 49.23 -6.63 -2.78
Watermark-Probe-2 (w. prompt 4.3)
Qwen2.5-1.5B -5.06 34.55 55.97 16.39 8.79 19.45 14.28 10.73 1840.44
OPT-2.7B -1.95 42.59 67.93 15.17 4.61 32.40 11.04 26.44 1073.13
Llama-3.2-3B -12.42 29.91 96.14 50.00 19.91 80.04 67.07 34.99 7702.12
Qwen2.5-3B -3.48 6.35 108.44 11.29 25.92 39.88 18.84 18.06 8209.12
Llama2-7B -2.87 24.87 40.05 32.68 15.62 35.50 25.11 47.74 6885.04
Mixtral-7B -0.87 6.09 54.39 11.14 13.49 49.02 111.12 14.83 1812.12
Qwen2.5-7B -2.48 8.90 185.88 10.50 13.06 7.64 12.40 13.04 1982.74
Llama-3.1-8B -64.31 25.24 104.77 10.03 49.23 38.47 81.36 31.35 12701.95
Llama2-13B -3.98 20.72 38.26 10.36 16.60 58.10 83.27 47.74 333.37
Table 8: Identification p-value for various LLMs with different watermarks and without watermarks, using our two methods: Watermark-Probe-v1 and Watermark-Probe-v2. represents high-confidence watermark and represents low-confidence watermark, while no color means no watermark. This table provides supplementary information on the similarity content in Table 1.
LLM N-Gram Fixed-Key-Set
Non KGW Aar KGW-Min KGW-Skip DiPmark $\gamma$-reweight EXP-Edit ITS-Edit
Watermark-Probe-v1 (w. prompt 4.3)
Qwen2.5-1.5B 1 2.9e-410 5.9e-3319 3.2e-49 1.3e-130 3.6e-244 2.0e-681 1 1
OPT-2.7B 1 1.1e-528 3.4e-3024 7.7e-63 1.1e-280 2.6e-1891 2.9e-545 1-9.3e-2 1-1.0e-1
Llama-3.2-3B 1 1.6e-532 1.4e-1369 1.3e-1425 5.2e-1101 4.4e-1653 7.9e-1269 1 1
Qwen2.5-3B 1 2.7e-507 1.8e-3509 2.1e-49 8.3e-84925 1.7e-687 1.6e-403 1 1-2.4e-2
Llama2-7B 1 3.6e-198 4.3e-2583 7.6e-144 8.7e-195 7.0e-430 4.4e-2460 1-1.1e-3 1
Mixtral-7B 1 1.4e-138 8.0e-354 1.7e-276 3.4e-89 8.1e-289 3.9e-4102 1 1
Qwen2.5-7B 1-1.2e-1 2.1e-320 2.0e-198 7.3e-116 4.2e-553 7.9e-489 3.6e-556 1 1
Llama-3.1-8B 1 8.4e-93 4.4e-4476 7.7e-186 3.9e-175 2.5e-185 3.6e-3421 1 1
Llama2-13B 1 4.0e-188 1.1e-1372 4.0e-32 8.1e-73 2.7e-203 4.3e-529 1 1-2.7e-3
Watermark-Probe-v2 (w. prompt 4.3)
Qwen2.5-1.5B 1 7.1e-262 4.1e-683 1.1e-60 7.5e-19 1.5e-84 1.5e-46 3.7e-27 9.8e-735530
OPT-2.7B 1-2.6e-2 1.2e-396 5.6e-1005 2.8e-52 2.0e-6 1.4e-230 1.2e-28 2.4e-154 1.2e-250072
Llama-3.2-3B 1 7.3e-197 3.5e-2010 1.1e-545 1.7e-88 3.7e-1394 9.2e-980 1.6e-268 2.5e-12881755
Qwen2.5-3B 1 1.1e-10 1.2e-2556 7.4e-30 2.0e-148 4.4e-348 1.8e-79 3.3e-73 7.3e-14633482
Llama2-7B 1-2.1e-3 7.9e-137 4.9e-351 1.5e-234 2.7e-55 2.5e-276 1.9e-139 1.0e-497 4.3e-10293604
Mixtral-7B 1-1.9e-1 5.6e-10 3.1e-645 4.0e-29 9.0e-42 1.3e-524 2.0e-2684 4.7e-50 6.5e-713068
Qwen2.5-7B 1-6.6e-3 2.8e-19 3.9e-7506 4.3e-26 2.8e-39 1.1e-14 1.3e-35 3.6e-39 3.1e-853666
Llama-3.1-8B 1 7.3e-141 1.0e-2386 5.6e-24 4.3e-529 4.5e-324 2.0e-1440 4.9e-216 7.5e-35034440
Llama2-13B 1 1.1e-95 1.4e-320 1.9e-25 3.5e-62 6.8e-736 1.0e-1508 1.0e-497 2.0e-24136
Appendix F Details of Tested Watermarking Algorithms
To help understand the watermarking algorithms related to the experiments in this paper, we provide detailed information for all watermarking algorithms in Table 5, including their names, references, types, and brief descriptions. All our experiments were conducted using the MarkLLM (Pan et al., 2024a) framework.
Appendix G Ablation of Rank Transformation
To illustrate the importance of the rank transformation mentioned in Equation 9, we present in Figure 5 the variation of z-scores at different temperatures without using rank transformation. It can be observed that without rank transformation, the z-scores for Unwatermarked LLMs are significantly higher, especially at lower temperatures. Comparing the left plots in Figures 3 and 5, we can see that rank transformation effectively reduces the z-scores of Unwatermarked LLMs, making the identification and detection more stable.
Appendix H Supplementary Z-Scores and P-values
To facilitate a better understanding of the statistical methods used in identifying watermarked LLMs, we provide detailed information including z-scores and p-values for Table 1 and Table 2 in this section.
Specifically, Table 6 provides supplementary z-score information for Table 2, Table 7 provides supplementary z-score information for Table 1, and Table 8 provides supplementary p-value information for Table 1.
For all experiments, we consider a z-score below 4 to indicate no watermark, between 4 and 10 to indicate a watermark with relatively low confidence, and above 10 to indicate a watermark with high confidence.